Enter a numerical value in the blank. 3 The ion PO4-has valence electrons.
Answer #1

unun Answerg Given in pol Valency ēns = 5+4(6) +3 =5+24+3 =32 ens present Note 8- D p valency ens=5 2) o Nalency ens z6